### v2

# A Survey on Hallucination in Large Vision-Language Models

Abstract
Recent development of Large Vision-Language Models (LVLMs) has attracted growing attention within the AI landscape for its practical implementation potential. However, “hallucination”, or more specifically, the misalignment between factual visual content and corresponding textual generation, poses a significant challenge of utilizing LVLMs. In this comprehensive survey, we dissect LVLM-related hallucinations in an attempt to establish an overview and facilitate future mitigation. Our scrutiny starts with a clarification of the concept of hallucinations in LVLMs, presenting a variety of hallucination symptoms and highlighting the unique challenges inherent in LVLM hallucinations. Subsequently, we outline the benchmarks and methodologies tailored specifically for evaluating hallucinations unique to LVLMs. Additionally, we delve into an investigation of the root causes of these hallucinations, encompassing insights from the training data and model components. We also critically review existing methods for mitigating hallucinations. The open questions and future directions pertaining to hallucinations within LVLMs are discussed to conclude this survey. To keep track of this field and continuously update our survey, we maintain a repository of relevant references at https://github.com/lhanchao777/LVLM-Hallucinations-Survey.
1 Introduction
In the rapidly evolving realm of artificial intelligence, large language models (LLMs) such as GPT-4 OpenAI (2023), LLaMA Touvron et al. (2023a), and LLaMA2 Touvron et al. (2023b) have made remarkable strides in natural language understanding (NLU) and generation (NLG). To harness the NLU and NLG capabilities of LLMs for vision-language tasks, a prevalent approach is inserting visual features as supplementary inputs to LLMs and aligning them with textual features. This method has been adapted in several large vision-language models (LVLMs) such as MiniGPT-4 Zhu et al. (2023), LLaVA Liu et al. (2023c), and LLaVA-1.5 Liu et al. (2023b). Despite the promising results demonstrated by existing LVLMs, an unignorable issue has been impeding their practical application: hallucination. Hallucination in LVLM denotes a disagreement between factual content of images and corresponding generated textual content, akin to the solely textual hallucination encountered in large language models Huang et al. (2023a).
Figure 1: Hallucination examples in LVLMs. Hallucination symptoms may manifest as deficiencies in various vision-language tasks like judgment and description, or factual errors in different visual semantics, such as objects, attributes, and relations.
Existing studies Rohrbach et al. (2018); Li et al. (2023b); Hu et al. (2023); Zhai et al. (2023) have addressed the issue of hallucination in image captioning models with a predominant focus on the “presence of objects”, specifically whether the objects depicted in a given image are accurately described by the text generated by the model. In contrast to image captioning models trained within closed domains, LVLMs leverage the robust understanding and expressive capabilities of LLMs to obtain more detailed and fluent generated descriptions. However, these enhanced abilities also diversify and potentially exacerbate hallucination, which is not only limited to object existence, but also manifested in the descriptive errors such as attribute and relation errors. We focus on visual hallucinations, referring to all inconsistencies between the semantic content conveyed by the image and the textual content generated by the model.
Hallucination symptoms in LVLMs are multifaceted. From a cognitive perspective, hallucinations can manifest as errors in true/false judgments and inaccuracies in the description of visual information. For instance, as illustrated in the first example of Figure 1, the model ill responds to questions such as “Is there a cat in the image?” and “Are there four birds in the image?”, demonstrating flawed factual discernment. Additionally, the second example shows an misalignment of the generated description and the visual facts. Meanwhile, a visual semantics perspective provides us with a ternary taxonomy: hallucination on object, attribute, and relationship. For example, the model generates non-existent objects in the image such as “laptop” and “small dog”, provides incorrect attribute descriptions such as describing the man being “long-haired”, and makes inaccurate assertions about the relationships between objects such as claiming the bicycle is “in front of” the man. Current methods evaluate these hallucination in LVLMs based on the models’ cognitive performance, focusing on two main aspect: non-hallucinatory generation and hallucination discrimination. The former involves a detailed analysis of the hallucinatory elements in the model’s response and quantifying their proportion. The latter, on the other hand, only requires a binary judgment of whether the response comprises any hallucinatory content. These approaches are comprehensively discussed in §3.
While the causes of hallucinations in LLMs Huang et al. (2023a) have been discussed extensively, LVLMs’ visual modality introduces unique challenges in analyzing these occurrences. We conduct a thorough analysis of hallucinations in LVLMs, focusing on the training data and model characteristics. Our analysis indicates that hallucinations in LVLMs are caused not only by the generative nature of LLMs but also by biased training data, the inability of vision encoders to accurately ground images, misalignment among different modalities, insufficient context attention, and many other factors. Following this, we provide a comprehensive overview of the existing hallucination mitigation methods. Corresponding to the causes, current mitigation approaches predominantly focus on the optimization of training data, refinement of various modules within LVLMs, and post-processing of generated outputs. These methods are employed to diminish the occurrence of hallucinations, thereby yielding more faithful responses. In the end, we list several considerable directions for developing the research of hallucinations in LVLMs.
In summary, this study aims to provide insights for the development of LVLMs and explore the opportunities and challenges related to LVLM hallucinations. This exploration not only helps us understand the limitations of current LVLMs, but also offers important guidance for future research and the development of more reliable and efficient LVLMs.
2 Hallucination in the Era of LVLM
2.1 Large Vision-Language Models
LVLMs are advanced multimodal models that process visual and textual data to solve compound tasks involving both vision and natural language. Incorporating LLMs’ capabilities, LVLMs are an evolution of previous Vision-Language Pre-trained Models (VLPMs) Long et al. (2022).
LVLM architectures typically comprise three components: a visual encoder, a modality connection module, and a LLM. The visual encoder, often an adaptation of the CLIP vision encoder Radford et al. (2021), transforms input images into visual tokens. The connection module is designed to align visual tokens with the word embedding space of the LLM, ensuring that the LLM can handle the visual information. There are various methods for modality alignment, including cross-attention Alayrac et al. (2022), adapters Gao et al. (2023), Q-Formers Li et al. (2023a); Dai et al. (2023a); Zhu et al. (2023), and simpler structures such as linear layers or multi-layer perceptrons (MLP) Liu et al. (2023c); Chen et al. (2023b); Liu et al. (2023b). The LLM functions like the central processing unit within LVLMs, receiving aligned visual and textual information and subsequently synthesizing this information to produce responses.
The training of LVLMs involves two key stages: (1) pre-training, where LVLMs acquire vision-language knowledge from aligned image-text pairs, and (2) instruction-tuning, during which LVLMs learn to follow human instructions using a varied task dataset. Following these stages, LVLMs can efficiently process and interpret visual and textual data, enabling them to reason about concepts in compound multimodal tasks like Visual Question Answering (VQA).
2.2 Hallucination in LVLMs
Hallucination in LVLMs refer to contradictions between the visual input (taken as ‘fact’) and the textual output of a LVLM. Through the lens of visual-language tasks, LVLM hallucination symptoms can be interpreted as deficiencies in judgement or description.
A judgement hallucination occurs when the model’s response to a user’s query or statement is in disagreement with the actual visual data. For example, as shown in Figure 1, when being presented with an image depicting three birds and inquired whether there is a cat in the picture, the model wrongly affirms with a “Yes”. A description hallucination, on the other hand, is a failure to faithfully depict the visual information. As exemplified in lower part of Figure 1, where the model inaccurately describes the man’s hair, the quantity and color of cups, the bicycle’s location, and fabricates non-existent objects such as laptop and dog.
From a semantic perspective, such misalignment can be characterized by claiming nonexistent objects, incorrect object attributes, or inaccurate object relations, as highlighted in different colors.
2.3 Unique Challenges regarding Hallucination in LVLMs
LVLMs handle vision-language tasks by incorporating both visual and language modules. However, the integration also pose unique challenges in hallucination detection, causal inference, and mitigation methods.
Hallucination Detection Difficulties
The multimodal nature of LVLM impedes detecting hallucinations. LVLM Hallucinations may manifest across various semantic dimensions, including, but not limited to the object, attribute, and relation Zhai et al. (2023); You et al. (2023). To detect these hallucinations comprehensively, the model is tasked not only on natural language understanding but also employing fine-grained visual annotations and aligning them precisely with the generated text.
Intertwined Causes
The causes of hallucinations in LVLMs are often multifaceted. On one hand, there are data-related issues shared by both LLMs and LVLMs Hu et al. (2023) such as misinformation, biases, and knowledge boundary limits. However, LVLMs uniquely suffer from their incorporation of visual data. For instance, visual uncertainty, such as unclear or distorted images, can exacerbate the language priors and statistical biases in LVLMs, leading to severer hallucinations Liu et al. (2023a).
Compound Mitigation Methods
Besides adapting LLM-targeted hallucination mitigation methods, such as data quality enhancement, encoding optimization, and aligning with human preferences, LVLM-specific methods also include refining visual representations and improving the multi-modal alignment. For instance, enlarging visual resolution has been suggested to effectively reduce hallucinations Bai et al. (2023). Nonetheless, training high-resolution visual encoders with extensive data can be highly resource-demanding. Therefore, it is imperative to investigate more cost-effective strategies for augmenting visual representations. Moreover, the significant gap between visual and textual tokens suggest that improving the vision-language token alignment can potentially lower the incidence of hallucinations Jiang et al. (2023).
3 Evaluation Methods and Benchmarks
{forest}
[…]
Evaluation Methods, ver [ Evaluation on
Non-Hallucinatory Generation [ Handcrafted
Pipeline Methods [ CHAIR Rohrbach et al. (2018),
CCEval Zhai et al. (2023),
FAITHSCORE Jing et al. (2023) , leaf, text width=15em ] ] [ Model-based
End-to-End Methods [ LLM-based Evaluation [ GAVIE Liu et al. (2023a),
MMHal-Bench Sun et al. (2023) , leaf, text width=15em ] ] [ Hallucination Data
Driven Model Evaluation [ M-HalDetect Gunjal et al. (2023),
HaELM Wang et al. (2023c) , leaf, text width=15em ] ] ] ] [ Evaluation on
Hallucination Discrimination [ POPE Li et al. (2023b),
CIEM Hu et al. (2023),
NOPE Lovenia et al. (2023) , leaf, text width=15em ] ] ]
Figure 2: Taxonomy of Hallucination Evaluation Methods for LVLMs
Table 1: Hallucination Evaluation Benchmarks for LVLMs. ‘Dis’ and ‘Gen’ stand for ‘Discriminative’ and ‘Generative’, respectively.
Benchmark Evaluation Size Hallucination Symptoms Metrics
POPE Li et al. (2023b) Dis 3000 Object Accuracy
NOPE Lovenia et al. (2023) Dis 17983 Object Accuracy
CIEM Hu et al. (2023) Dis 72941 Object Accuracy
M-HalDetect Gunjal et al. (2023) Gen 4000 Object & Attribute & Relation Reward Model Score
GAVIE Liu et al. (2023a) Gen 1000 Object & Attribute Accuracy & Relevancy
FAITHScore Jing et al. (2023) Gen 2000 Object & Attribute & Relation FAITHScore
HaELM Wang et al. (2023c) Gen 5000 Not Explicitly Stated Accuracy
MMHal-Bench Sun et al. (2023) Gen 96 Object & Attribute & Relation Rating Score
AMBER Wang et al. (2023b) Dis & Gen 15202 Object & Attribute & Relation AMBER Score
Establishing the concepts of hallucination in LVLM, we then turn to examine existing LVLM hallucination evaluation methods and benchmarks. Corresponding to the hallucination symptoms in description and judgement tasks mentioned in Figure 1, current evaluation methods can be classified into two major approaches: (1) assessing the model’s ability of non-hallucinatory content generation, and (2) evaluating the model’s ability of hallucination discrimination, as shown in Figure 2. Similarly, the benchmarks can also be categorized into discriminative and generative ones on the basis of evaluation tasks, as demonstrated in Table 1.
3.1 Evaluation on Non-Hallucinatory Generation
Evaluating non-hallucinatory generation is to measure the proportion of hallucinated content in the outputs. Currently, there are primarily two main types: handcrafted pipeline and model-based end-to-end methods.
Handcrafted Pipeline Methods
Handcrafted pipeline methods feature strong interpretability by manually designing multiple steps with specific and clear objectives. CHAIR Rohrbach et al. (2018) targets evaluating object hallucinations in image captioning by quantifying differences of objects between model generation and ground-truth captions. However, such conventional methods struggle with LVLMs’ vast object categories Lu et al. (2018). To address this, CCEval Zhai et al. (2023) employs object alignment module powered by GPT-4 before applying CHAIR. FAITHSCORE Jing et al. (2023) offers a reference-free and fine-grained evaluation method for evaluating the faithfulness of free-form model responses, which comprises identifying descriptive sub-sentences, extracting atomic facts, and comparing extracted facts against the corresponding input image.
Model-based End-to-End Methods
End-to-end methods directly assess the performance of LVLMs by evaluating their responses, such as scoring the responses or determining the presence of hallucinations in responses. Existing end-to-end evaluation methods can broadly be categorized into two types. The first type is LLM-based evaluation which implements an advanced LLM (e.g., GPT-4) to rate LVLM generated content based on hallucination Liu et al. (2023a); Sun et al. (2023). These methods leverage the LLM’s robust natural language understanding and processing capabilities. By integrating visual information (such as dense captions and object bounding boxes), user instructions, and model responses as input, LLMs are prompted to evaluate and score these responses. The second type is hallucination data driven model evaluation which constructs labelled hallucination datasets for fine-tuning models to detect hallucinations. Gunjal et al. (2023) creates the M-HalDetect dataset with annotated LVLM image descriptions and fine-tune an InstructBLIP Dai et al. (2023a) model on the dataset for hallucination identification. Similarly, Wang et al. (2023c) compiles a dataset of hallucinatory responses generated by ChatGPT and then fine-tune a LLaMA model with LoRA Hu et al. (2022) on this dataset to differentiate hallucinatory responses from non-hallucinatory ones.
3.2 Evaluation on Hallucination Discrimination
Hallucination discrimination evaluation approach aim to assess the hallucination discrimination ability of LVLMs. The methods follow this approach typically adopt a question-answering format, posing inquiries to LVLMs consisting of descriptions that agree or conflict with the content of the provided input image and assessing the models’ responses Li et al. (2023b); Hu et al. (2023); Lovenia et al. (2023). POPE Li et al. (2023b) designs binary (Yes-or-No) questions about object presence in images such as “Is there a person in the image?” to evaluate the hallucination discrimination ability of LVLMs. The objects asked in questions are selected under three distinct sampling strategies: random (selecting random absent objects), popular (choosing most frequent objects in dataset but absent in current image), and adversarial (selecting absent objects often co-occurring with present ones). CIEM Hu et al. (2023) is akin to POPE, yet it automates object selection by prompting ChatGPT. NOPE Lovenia et al. (2023) is another VQA-based method that is designed to evaluate LVLMs’ ability of discerning the absence of objects in visual queries, with correct responses being negative statements.
Figure 3: The causes and mitigation methods of hallucinations in LVLMs.
3.3 Evaluation Benchmarks
Distinct from standard LVLM benchmarks that assess general LVLM capabilities, LVLM hallucination benchmarks specifically target non-hallucinatory generation or hallucination discrimination. Many such benchmarks have been proposed along with the evaluation methods. Table 1 shows a synthesis of representative benchmarks. These benchmarks are categorized based on the type of evaluation approach: Discriminative (Dis) or Generative (Gen).
Discriminative Benchmarks
POPE Li et al. (2023b), NOPE Lovenia et al. (2023), and CIEM Hu et al. (2023) are discriminative benchmarks, and their dataset sizes are 3000, 17983, and 72941, respectively. All of these three benchmarks focus only on object hallucinations and adopt accuracy as the evaluation metric, which is obtained by inquiring whether objects are present in images and comparing model responses with ground-truth answers.
Generative Benchmarks
As delineated in Table 1, current research emphasizes generative benchmarks over discriminative ones. While discriminative benchmarks primarily evaluate hallucinations at the object level, generative benchmarks extend their scope to include a broader spectrum of hallucinations, including attribute and relation ones Gunjal et al. (2023); Liu et al. (2023a); Jing et al. (2023); Wang et al. (2023c); Sun et al. (2023). Notably, AMBER Wang et al. (2023b) stands out as a comprehensive benchmark that integrates both generative and discriminative tasks. Furthermore, we observe that the metrics employed in generative benchmarks tend to be more complex and varied, in contrast to the reliance on accuracy metrics in discriminative benchmarks. This is probably because, distinct from discriminative methods that mostly compare model outputs with ground-truth answers, generative evaluation methods require tailored metrics to target specific hallucination categories in an attempt to analyze the responses generated by LVLMs.
4 Causes of LVLM Hallucinations
Hallucinations in LVLMs can arise from various factors. As depicted in Figure 3, in correspondence with LVLM pipeline. These causes can be categorized into the following aspects:
4.1 Hallucinations from Data
The quality of training data largely impacts training efficiency and model performance. However, existing LVLM training data suffer from several quality challenges, fostering certain types of hallucinations.
Data Bias
One notable issue of data bias is the presence of distribution imbalance in existing training data, particularly in factual judgment QA pairs where the majority of the answers are “Yes” Hu et al. (2023); Liu et al. (2023a). LVLMs trained on such biased data tend to answer “Yes” consistently and provide explanations even for incorrect or misleading prompts. Another issue is data homogeneity, which impedes the model’s ability to understand visual information and execute instructions under various environments You et al. (2023); Liu et al. (2023a). For example, MiniGPT-4 can only describe images irrespective of the specific question posed by users due to the lack of diversified instruction learning. Similarly, LLaVA struggles to accurately describe local visual relations due to the limited varieties in visual information it was trained on.
Annotation Irrelevance
A large amount of instruction data is primarily synthesized from image-caption and detection data using LLMs. Nevertheless, due to the unreliability of these generative models, this approach encounters a challenge of annotation relevance: the generated long instructions often contains objects, attributes, and relationships that may not correspond to the fine-grained content depicted in the images Liu et al. (2023a). Training on such data also catalyzes hallucinations.
4.2 Hallucinations from Vision Encoder
The widely adopted vision encoders in LVLMs are derived from CLIP, which maps visual and textual features to the same space by contrastive learning. Though CLIP achieves excellent performance on various visual understanding tasks, it is nonetheless limited in fully expressing visual information, such as restricted visual resolution and fine-grained visual semantics.
Limited Visual Resolution
Higher image resolution can make visual encoder more accurate in object recognition and perceive more visual details, thus alleviating hallucinations Zhai et al. (2023); Li et al. (2023c). However, handling a wider range of image resolutions is computation-wise demanding, hence existing models usually use smaller image resolution, such as $224\times 224224 × 224$ and $336\times 336336 × 336$ pixels for LLaVA and LLaVA-1.5 respectively.
Fine-grained Visual Semantics
CLIP aligns visual content and word embeddings through contrastive learning that mainly focuses only on salient objects, failing to capture fine-grained aspects of an image and resulting in hallucinations Jain et al. (2023); Cho et al. (2022), such as background description, object counting, and object relation.
4.3 Hallucinations from Modality Aligning
The connection module projects visual features into the LLM’s word embedding space, aligning visual and textual modalities. Consequently, misalignment could be a key factor in hallucination generation Sun et al. (2023); Zhao et al. (2023b). Furthermore, Jiang et al. (2023) reveal that even in advanced LVLMs like MiniGPT-4 and LLaVA, the gap between visual and textual features stands significant and lead to hallucinations. The simplicity of connection module and constraint of tokens are two main causes.
Connection Module Simplicity
Simple structures, such as linear layers, are commonly employed as connection modules to align visual and textual modalities. Although cost-efficient, the simplicity of the connection modules hinder comprehensive multimodal connection Liu et al. (2023b), thereby increasing the risk of hallucinations.
Limited Token Constraints
Q-Former is a widely used multimodal alignment module in LVLMs such as BLIP-2, InstructBLIP and MiniGPT-4. It is trained with various objectives to encode a predetermined number of random initialized tokens (e.g., 32) into visual features that are aligned with text. However, the restricted quantity of these tokens prevents encoding all the information present in images during the alignment process Yin et al. (2023a); Chen et al. (2023c), and consequent information loss raises the risk of hallucinations.
4.4 Hallucinations from LLM
LLMs are pivotal in the architecture of LVLMs, as LLMs significantly enhance LVLMs ability to process complex multimodal tasks. However, this integration also introduce the inherent hallucination challenges of LLMs to LVLMs. A growing body of literature has explored various factors contributing to hallucinations in LLMs Ji et al. (2023); Zhang et al. (2023b); Huang et al. (2023a). In this section, we briefly summarize potential causes of hallucinations originating from LLMs within the scope of vision-language.
Insufficient Context Attention
Insufficient context attention is the phenomenon where the model, during the decoding process, focuses only on partial information from the context. The phenomenon includes focusing excessively on the current segment of generated content while ignoring input visual information Wang et al. (2023a); Lee et al. (2023), prioritizing language patterns to produce fluent yet inaccurate content Wang et al. (2023c), and concentrating on partial summarizing tokens of the generated content, resulting in fabrication and hallucinatory content Huang et al. (2023b).
Stochastic Sampling Decoding
Stochastic sampling introduces randomness into the decoding process and is a commonly used decoding strategy in LLMs. It helps prevent the generation of low-quality text often associated with sequences of high likelihood and serves to enrich the generated content Holtzman et al. (2020). However, the incorporation of randomness in the model also amplify the risk of hallucinations Chuang et al. (2023); Dziri et al. (2021).
Capability Misalignment
Capability misalignment in LLMs is the disparity between the model’s inherent capabilities that established during the pre-training phase, and the expanded requirements imposed during the instruction tuning. This misalignment leads LLMs to produce responses beyond their established knowledge limits, increasing the potential for hallucinatory content Huang et al. (2023a).
5 Mitigation of LVLM Hallucination
Various methods have been proposed to mitigate hallucinations in LVLMs. As depicted in Figure 3, these methods are tailored to address the identified causes. In this section, we elaborate on them from the following perspectives:
5.1 Mitigation for Data
Optimizing training data is a direct and effective way to alleviate hallucinations. Regarding the explored deficiency in current datasets (§4.1), some measures are taken to address the issues in the following two perspectives:
Bias Mitigation
Considering the imbalance of positive and negative samples in the training data, CIEM Hu et al. (2023) utilizes off-shelf LLMs to generate contrastive question-answer pairs from annotated image-text datasets. These pairs are then leveraged in introduced Contrastive Instruction Tuning (CIT). LRV-Instruction Liu et al. (2023a) proposes a large and diverse dataset consisting of 400,000 visual instructions. This dataset not only includes positive visual instructions but also incorporates negative ones on three different semantic levels. Furthermore, delving into a fine-grained manner, Ferret You et al. (2023) proposes to mine 95,000 negative samples by replacing the original categories, attributes, or quantity information with analogous fake ones. This method effectively enhances the model robustness.
Annotation Enrichment
Constructing richly annotated datasets helps mitigating hallucination by supervising LVLMs to extract visual content accurately and align modalities comprehensively. One notable dataset is M-HalDetect Gunjal et al. (2023). It consists of 4,000 image-description pairs sourced from COCO labelled with object existence, relative positions, and attributes. Another valuable dataset is GRIT You et al. (2023). GRIT features 1.1 million samples with refer-and-ground instructions over hierarchical spatial knowledge. In addition, Zhai et al. (2023) develops two datasets: one with RAM-system Zhang et al. (2023a) grounded objects and the other with objects grounded by RAM or generated by LLMs. Lastly, Lu et al. (2023) generated 180,000 fine-grained instruction samples, covering multi-round conversations, vision-prompted recognition, and fact-checking.
5.2 Mitigation for Vision Encoder
Scaling-up Vision Resolution
Bai et al. (2023) have shown the effectiveness of gradual enlargement of image resolution from $224\times 224224 × 224$ to $448\times 448448 × 448$. MONKEY Li et al. (2023c) processes high-resolution images by dividing them into patches with the size matched the vision encoder, and local features extracted from patches are processed by the LLM. Chen et al. (2023c) proposes InternVL that scales up the vision encoder to 6 billion parameters and can process images with widths ranging from 1,664 to 6,144 pixels. However, this approach is resource-demanding for pretraining with large-scale data.
Perceptual Enhancement
Most existing LVLMs employ the ViT from CLIP as the vision encoder, which mainly focuses only on salient objects and inevitably omit some visual cues. To enhance the object-level perceptual capability, Jain et al. (2023) proposes VCoder that uses extra perception modalities, such as segmentation map, depth map, or both, as the control inputs through additional vision encoders. To enhance the spatial awareness capability, Zhao et al. (2023a) suggests introducing addition pre-trained models to acquire spatial position information and scene graph details, which are then used to guide the LVLMs in addressing user queries.
5.3 Mitigation for Connection Module
Connection Module Enhancing
As discussed in §4.3, the limited capacity of the connection module hinders perfect modality alignment, increasing hallucination risks. To better align the vision and language modalities, researchers have developed more capable connection modules recently. For instance, LLaVA-1.5 enhances the connection module in LLaVA by upgrading from a single linear layer to a MLP. Moreover, Chen et al. (2023c) utilizes LLaMA2 to construct QLLaMA, which significantly outperforms the Q-Former in aligning visual features with text.
Alignment Training Optimizing
Strengthening the alignment training process is an effective method to enhance modality alignment and reduce the risk of hallucinations. Jiang et al. (2023) observe a significant gap between visual and textual tokens in LVLMs. By explicitly adding new learning objectives to bring visual and textual tokens closer, the hallucinations can be reduced. On the other hand, Sun et al. (2023) employ Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2020) to align different modalities, achieving hallucination reduction.
5.4 Mitigation for LLM
Decoding Optimization
In response to insufficient context attention, some recent research attempts to enable the model to focus on proper contexts during decoding, thereby mitigating hallucinations. Huang et al. (2023b) notice LVLMs may generate hallucinations by over-focusing on few summary tokens and neglecting image tokens. To mitigate hallucinations, they develop a novel decoding strategy named OPERA that modifies the beam search process with a weighted scoring system to de-prioritize over-trusted candidates and a retrospection mechanism for re-evaluating decision points. Leng et al. (2023) focused on visual uncertainty and its impact on hallucinations caused by statistical bias and language priors. They proposed a visual contrastive decoding strategy, which contrasts outputs from original and altered visuals, helping correct the model’s over-reliance on language priors and statistical biases.
Aligning with Human
Training LVLMs to align with human preferences can improve the quality of model responses. Recently, some efforts focus on preference-aligned training, where non-hallucinatory responses are favored, while hallucinatory contents are not preferred. LLaVA-RLHF Sun et al. (2023) pioneers the use of RLHF for training LVLMs to align with human preferences. The authors first train a reward model to mirror human preferences. Subsequently, they apply a fact augmented RLHF and the reward model to align the LVLMs with human preferences. Another category of methods Gunjal et al. (2023); Zhao et al. (2023b); Yu et al. (2023) are based on Direct Preference Optimization (DPO) Rafailov et al. (2023) that trains LLMs directly from human preference data, avoiding the complexities of reward-model-based reinforcement learning methods like RLHF. After constructing hallucinatory labelled training data, these work present DPO variants, notably HA-DPO Zhao et al. (2023b) and FDPO Gunjal et al. (2023), to train the LVLMs better align with human and generate non-hallucinatory responses.
5.5 Mitigation via Post-processing
Beyond manipulating modules of LVLMs, hallucinations can be mitigated through post-processing or output editing via an additional module or operations. The post-processing methods input visual data, user instructions, and LVLM responses, and outputs hallucination-free, refined responses.
More recently, two post-processing methods were introduced to reduce hallucinations in generated image descriptions: LURE Zhou et al. (2023) and Woodpecker Yin et al. (2023b). LURE leverages insights into the causative factors of object hallucination to train a revisor model based on MiniGPT-4. Woodpecker is a training-free post-processing method comprising five stages that extract key concepts, formulate questions, apply visual knowledge validation, generate a structured visual knowledge base, and refine the generated content by correcting inaccuracies and integrating evidence from the knowledge base.
6 Future Directions
The research on hallucinations in LVLMs is still in its early stage, leaving considerable room for development and improvement. We detail these areas of further research below:
Supervision Objective
Meticulous supervision objectives are crucial for models to acquire desired capabilities or suppress unsought symptoms. Current LVLMs predominantly employ coarse-grained (e.g., image-caption) Li et al. (2023a) alignments and whole-image instruction tuning, which limits their ability to understand complex visual details. More detailed, spatially-aware supervision such as object masked language modeling Dai et al. (2023b) for alignment and position-enhanced instruction following Chen et al. (2023a) hold immense potentials for further advancements.
Enriching Modalities
Integrating multi-modal signals not only expands the capabilities of models but also notably enhances their performance on single-modal tasks Girdhar et al. (2023). This can be attributed to high-quality representation learning from correlated and complementary information. Recent LVLMs research made strides in incorporating more modalities (e.g., videos) through methods including through cache retrieval Han et al. (2023) and pre-alignment Lin et al. (2023). These studies show benefits in addressing the “existence” issue. Moving forward, exploration into designing advanced paradigms for coupling multiple modalities is a worthwhile avenue of investigation.
LVLM as Agent
Empowered by LLMs, the understanding and reasoning capabilities of LVLMs are propelled to unprecedented levels. Yet, accurately perceiving and expressing specific visual details (like numbers in tables) is still challenging. Rather than using more advanced visual encoders with extensive training, a practical alternative is to teach LVLMs to utilize visual tools (e.g., detection and segmentation models) and reason with accurate and structured outputs. Recent research on “LVLMs as agent”  Liu et al. (2023d) demonstrates impressive fidelity in tasks like OCR-VQA, object tagging, and grounding, validating the viability of this direction.
Delving into Interpretability
The post-processing methods (§5.5) address hallucinations in LVLMs by refining generated outputs, but require careful design and higher computational costs. Further study is needed on LVLMs’ internal hallucination mechanisms and potential solutions. Several decoding strategies (§5.4) and model editing techniques Cheng et al. (2023) show good reliability and effectiveness against hallucination. Future in-depth research into LVLMs’ defects could lead to promising interpretability-oriented solutions.
7 Conclusion
Equipped with advanced vision encoders, robust LLMs, and modality connection modules, LVLMs excel in open-domain vision-language tasks. However, hallucinations significantly challenge LVLMs’ practical application. In this survey, we conduct a meticulous investigation into the phenomenon of hallucinations in LVLMs. This exploration introduces innovative evaluation methods along with pertinent benchmarks, provides a detailed analysis of the fundamental causes behind these hallucinations, and discusses effective mitigation approaches. We also delve into the existing challenges and discuss potential directions. This survey aims to lay a foundation for tackle complexities of hallucinations in LVLMs and facilitate future research towards practical implementation of these models in various applications.